OpenStudy (anonymous):

In a certain population of 540 cats, a single gene with two alleles controls thickness of the cats' fur. Allele C provides for thick fur and is dominant, and c provides thin fur. In this population, 49 cats have thin fur. What are the frequencies of the two alleles?

OpenStudy (anonymous):

we find the frequency of homozygous recessive individuals f(cc)=49540=0.0907≈0.091 From the expanded Hardy-weinberg equation we know that c2=f(cc)=0.091→c=0.091−−−−√=0.30 f(C)=1−0.30=0.70 Therefore, allele C is 0.7 and allele c is 0.3